LKD Co. has 10 percent coupon bonds with a YTM of 8.6 percent. The current yield on these bonds is 9.5 percent. How many years do these bonds have left until they mature?

Years until maturity
Answer #1

There are 5 components requires in bond calculation – price, face value, coupon interest, YTM, and number of years.

Here, the number of years is missing but all the other 4 are available as below:

Face value = F = $1,000 (assumed)

Coupon interest = C = $1,000 × 10% = $100

Price = V = Coupon interest / Current yield = $100 / 0.095 = $1,052.63

YTM = 0.086

By the formula:

V = (C/YTM) × [{(1 + YTM)^n – 1} / (1 + YTM)^n] + {F / (1 + YTM)^n}

1,052.63 = (100/0.086) × [{(1 + 0.086)^n – 1} / (1 + 0.086)^n] + {1,000 / (1 + 0.086)^n}

1,052.63 = 1,162.79 × [(1.086^n – 1) / 1.086^n] + (1,000 / 1.086^n)

Now by the use of solver, n = 5

Answer: the bond matures in 5 years.